OpenStudy (anonymous):

At a particular baseball stadium the outfield wall is 310 feet away from home plate and the wall is 6 feet tall. A player hit the ball when it was 3 feet off of the ground with an initial upward velocity of 64 feet per second and an initial outward velocity of 77 feet per second. Write the function h(x) that gives the height of the ball after time (t).

OpenStudy (anonymous):

i am really rusty at this - i know i could do these at one time. I'll Be back in half hour with the answer if u still waiting.

OpenStudy (anonymous):

O.k cool thanks! Anyone else know how to do this?

OpenStudy (anonymous):

ok you use one of the constant acceleration equation take the acceleration due to gravity to be 32 ft/s/s h(x) = ut + 0.5 g t^2 where u = initial upward velocity, t=time and g = acceleration due gravity so h(x) = 64t - 0.5*32t^2 h(x) = 64t - 16t^2 the negative is due to gravity acting in opposite direction to motion of the ball oh sorry - i missed out the initial 3 ft so its: h(x) = 64t - 16t^2 + 3

OpenStudy (anonymous):

Thank you so much! So based off of this, how long does the ball stay in the air?

OpenStudy (anonymous):

ok u can use the same equation to work this now h = 0 so 0 = 64t - 16t^2 + 3 solving this t use the formula: t = -64 + sqrt(64^2 - 4 * -16 * 3 / -32 this gives t = 4.05

OpenStudy (anonymous):

seconds that is

OpenStudy (anonymous):

Wow. You and amistre64 were really close! He got 4.04 seconds. I understand your explanation though. Thanks again for the help jimmyrep!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

plz click good answer to close question when u are done

OpenStudy (anonymous):

Oh wait! Another thought: How far will the ball travel? This is based off of the time (4.04 seconds).

OpenStudy (anonymous):

How long did the ball take to return to a height of 6 feet?

OpenStudy (anonymous):

AND FINALLY: How far did the ball travel before it returned to a height of 6 feet?

OpenStudy (anonymous):

ok - i'// take 4.04 seconds - igot 4.046 and rounded it up - but whatever the horiznontal distance the ball will travel is given by v = distance / time where v = initial outward velocity distance = v * t = 77 * 4.04 = 311 feet

OpenStudy (anonymous):

Hmm. O.k got it. Now I just need help on the last two questions.

OpenStudy (amistre64):

the distance the ball travels is the outward velocity times 4.04 right?

OpenStudy (anonymous):

it started at 3 ft so at 6ft it will be 3 ft higher and so we can use 3 = 64t - 16t^2 solving this the time is 3.95 seconds

OpenStudy (amistre64):

77 ft per sec times 4.04 secs = distance traversed

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Wait so how far will the ball travel?

OpenStudy (amistre64):

it lands abt 311 feet away; how long is the path that it takes thru the air tho?

OpenStudy (anonymous):

distance travelled when at height 6ft = 3.95 * 77 = 304.2 ft

OpenStudy (anonymous):

How long did the ball take to return to a height of 6 feet?

OpenStudy (anonymous):

distance through the air? that will be a parabola right?

OpenStudy (amistre64):

"How far did the ball travel before it returned to a height of 6 feet?" I interpret this to mean that the ball is on the downward arc ....

OpenStudy (amistre64):

yes, the distance along the parabola; but I dont think that was the right interp for the question :)

OpenStudy (anonymous):

right - thats what I thought

OpenStudy (amistre64):

(-16t^2 +64t +3) - 6 = 0 tells us the times at h=6ft right?

OpenStudy (anonymous):

yes - i think thats right

OpenStudy (amistre64):

i get the upward part is abt; .05 secs the downward is abt; 3.95 secs

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

How long did the ball take to return to a height of 6 feet? The answer to this would be which one though?

OpenStudy (amistre64):

to return means to come back to; it can only come back to 6ft if its coming back down again; which is the farthest one away and the longest time to get there

OpenStudy (anonymous):

Oh so downward

OpenStudy (anonymous):

So now back to distance. How far did the ball travel before it returned to a height of 6 feet?

OpenStudy (amistre64):

jimmyrep said: distance travelled when at height 6ft = 3.95 * 77 = 304.2 ft

OpenStudy (anonymous):

i'm struggling a bit with my memory of doing these the horizontal component of the flight is constant - no acceleration so if the time to get to 6ft above the ground is 3.95 sec then distance = time x speed = 3.95 x 77 = 304.2 ft i'm about 95 % sure thats correct

OpenStudy (anonymous):

Are either of you good at tessellations?

OpenStudy (amistre64):

it correct; we assume that there is no loss of acceleration; or that the loss is neglible given the short amount of time

OpenStudy (anonymous):

So this answer is the same answer for the question of How far will the ball travel?

OpenStudy (amistre64):

4.04 * 77 is not the same answer as: 3.95 * 77 .....

OpenStudy (amistre64):

one measure total distance traveled while airborne; the other measures total distance traveled as it arcs thru 6 ft

OpenStudy (anonymous):

So 3.95*77 is how we would get the answer to how far the ball traveled?

OpenStudy (anonymous):

Which would be 304.15 feet?

OpenStudy (anonymous):

yes - thats the horizontal distance travelled

OpenStudy (anonymous):

O.k and the ball traveled 304.2 feet before it returned to a height of 6 feet? Just making sure.

OpenStudy (anonymous):

well i think that this distance is the horizontal distance travelled from the player to a point directly below the ball when it is descending at a height of 6 feet. Thats how I have interpreted the question.

OpenStudy (anonymous):

Hmm. O.k. Well THANK YOU SO MUCH GUYS FOR YOUR HELP! I couldn't have done it without YOU.

OpenStudy (anonymous):

Jimmy, are you familiar with tessellations?

OpenStudy (anonymous):

sorry mathkid i'm not

OpenStudy (anonymous):

its all good

OpenStudy (anonymous):

What was the maximum height the ball reached?

OpenStudy (anonymous):

are u still there ? i had to go away a while

OpenStudy (anonymous):

Yes, I just need to find out what was the maximum height the ball reached.

OpenStudy (anonymous):

Would it be 512 feet?

OpenStudy (anonymous):

v^2 = u^2 + 2 xg x h at maxm height the speed v = 0 0 = 64^2 - 2 *32 * h h = 4096 / 64 = 64 add initial 3 ft height = 67 ft

OpenStudy (anonymous):

Thanks! :D

OpenStudy (anonymous):

JIMMYREP, i posted another problem... if you wouldnt mind :/

OpenStudy (anonymous):

ok